given that y(x) =x^-1is one solution of the differential equation 2 x^2 y ′′ +3x y' − y = 0 x>0 find a second linearly independent solution of the equation
1
Expert's answer
2020-03-18T15:39:53-0400

Let's assume that "y_2=x^{-1}v(x)" is the second linearlyindependent solution of the equation.

Substituting it to the equation leads us to

"2x^2(v''x^{-1}-2v'x^{-2}+2vx^{-3})+3x(v'x^{-1}-vx^{-2})-vx^{-1}=0"

Since "x^{-1}" is the solution of the equation, we get


"2x^2(v''x^{-1}-2v'x^{-2})+3xv'x^{-1}=0"

"2xv''-4v'+3v'=0"

"v''-\frac{v'}{2x}=0"

Assume

"v'=w"

We get

"w'-\frac{w}{2x}=0"

"\frac{dw}{w}=\frac{dx}{2x}"

"w=\sqrt{x}+c_1"

Therefore

"v=\int(\sqrt{x}+c_1)dx=\frac{2}{3}x^{3/2}+c_1x+c_2"

Assuming "c_1=c_2=0" we get the second linearly independent solution in form "y_2=\frac{2}{3}x^{2/3}x^{-1}=\frac{2}{3}x^{-1/3}"